# Quasiparticle cooling algorithms for quantum many-body state preparation

### v1

@@ III.3 Finite density of auxiliaries
To sum up, the analysis above demonstrates efficiency of MCP with a finite density of auxiliaries for removing quasiparticles after $O(\theta^{-2})$ cooling cycles, in the PM phase. In agreement with the arguments of Ref. [43], we also find that topological, domain-wall type quasiparticles are more challenging to remove. Nevertheless, in an idealized system without noise, our cooling protocol allows preparation of the ground state with fidelity arbitrarily close to 1.
IV Noise effects
@@
Having established that the quasiparticle cooling algorithm can be used to efficiently prepare correlated ground states, we now turn to the effects of unwanted decoherence. Each experimental platform comes with its own sources of noise, and our goal will be to understand the limitations it imposes on the cooling efficiency. We will focus on the single-qubit decay and dephasing processes, relevant in systems of superconducting qubits. The kinetic equation can be generalized to include transitions mediated by weak noise. Using this modified kinetic theory, we analyze the dependence of the steady state quasiparticle population on the noise strength.
IV.1 Rate equation in the presence of noise
@@
Considering continuous time evolution, single-qubit decoherence processes can be conveniently described within Lindblad formalism [50]. The evolution of the system’s density matrix $\rho_{S}$ is given by (for the moment, we ignore presence of auxiliary qubits):
@@
and following Eqs. (16-23) we obtain the noisy rate equation, which includes cooling transitions $R_{\vec{\alpha}\vec{\beta}}$ and and an additional contribution $R_{\vec{\alpha}\vec{\beta}}^{\gamma}$ to transition rates from the noise terms:
@@
$R_{\vec{\alpha}\vec{\beta}}$ $=\theta^{2}\sum_{j=1}^{N_{A}}|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}|\bra{\vec{\beta}}\hat{\sigma}_{Q_{j}}^{+}\ket{\vec{\alpha}}|^{2},$ (52)
@@
Note that the noise is essentially ‘infinite temperature’ and has no term analogous to the spectral filtering function $F_{h,T}$.
IV.2 Application: cooling the transverse-field Ising model
Figure 5: Performance of the cooling algorithm under decoherence. (a) Quasiparticle density in the steady state, as a function of noise ratio $\gamma/\theta^{2}$, in PM (red) and AFM (blue) phases, for different system sizes $N_{S}$ from 10 to 80. We use MCP with fixed $T=2\beta=6\pi/\Delta$. At weak noise, PM quasiparticle density scales linearly with $\gamma$, while AFM displays a crossover from $n\propto\gamma$ to $n\propto\sqrt{\gamma}$ behaviour as noise strength is reduced. (b) Log-fidelity between steady state and quasiparticle vacuum, for MCP (solid lines) and SCP (dashed lines). We fix $T=2\beta$ and optimize over parameter $\beta$ for each protocol.
Next, we apply the noisy kinetic equation to cooling of a transverse-field Ising model in the presence of decay to the qubit ground state, and dephasing noise. This corresponds to $\mu=d,\phi$, with jump operators
@@ IV.2 Application: cooling the transverse-field Ising model
We assume homogeneous decoherence rates for simplicity, $\gamma_{j\phi}=\gamma_{\phi}$, $\gamma_{jd}=\gamma_{d}$, and rescale the dephasing rate, $\gamma_{\phi}\to\gamma_{\phi}/2$, following standard conventions in the literature on open quantum system.
@@
The matrix elements satisfy the selection rules discussed in Sec. III. We now summarise the main behaviour of the noisy rate equation in the two phases, with a more detailed discussion of matrix elements for $\hat{Z}_{j}$ and $\hat{\sigma}^{\pm}_{j}$ operators given in Appendix D. In the PM phase, single particle loss and gain processes occur throughout the chain (cooling and decay contributions); quasiparticle pair creation/annihilation and number-conserving scattering events occur due to dephasing. In the AFM phase, on account of the domain-wall nature of excitations, only two-particle scattering and pair creation/annihilation processes can occur in the bulk, and they are induced by cooling, decay and dephasing. Single particle transitions require an interaction with the Majorana edge mode, and are localized to the chain edge.
We have compared the predictions of the noisy kinetic equation and MPS simulations in the presence of noise, finding again a good agreement in the parameter regimes accessible to MPS (moderate value of noise strength, or moderate times at a very weak noise). We refer the reader to Appendix F for a comparison of the steady state occupations, also for different coupling strengths. The rate equation, however, allows us to study system sizes and parameter regimes far beyond the reach of matrix product state simulations  [^5] The steady state can be efficiently found by root-finding algorithms, after the cooling and noise-induced rates have been calculated via Wick’s theorem. In practice, the system reaches its steady state quickly and we evolve the rate equation in time until the quasiparticle occupations have effectively converged..
In Fig. 5a, we illustrate the quasiparticle density in the steady state, obtained from the kinetic equation, as a function of the noise strength. For simplicity, we choose $\gamma_{d}=\gamma_{\phi}=\gamma$ and fix auxiliary density $N_{A}=N_{S}$. We focus on the MCP with $T=2\beta=6\pi/\Delta$. We observe a qualitatively different behavior in the two phases, in line with above results on noiseless cooling. In the PM phase, the steady state quasiparticle density decays as a power-law function of noise, $n^{\infty}_{\rm PM}\propto\gamma^{a}$, $a\approx 1$, with little dependence on the system size. In the AFM phase, in contrast, the density displays a power-law scaling which depends on system size, $n^{\infty}_{\rm AFM}\propto\gamma^{b(N_{S})}$. For small systems and weak noise, we observe similar scaling to the PM, with $b\approx 1$; as either the system size or noise strength is increased, we find a crossover to a different power $b\approx 1/2$.
@@
where $C_{1},C_{\gamma}$ are constants of order one. We assumed that MCP is in the regime where heating processes due to coupling to auxiliaries are strongly suppressed, and neglected the dependence of $C_{1},C_{\gamma}$ on $k$. For the noise we neglect terms $O(\gamma Tn)$, since we are interested in the regime of weak noise and low density. Then, the expected quasiparticle density $n^{\infty}$ is given by
@@
(strictly speaking, this expression holds for small quasimomenta, $k\ll 1$, but we will see below that the steady state quasiparticle density is determined mostly by this quasimomentum range). The processes removing quasiparticle pairs can be modeled by a term $-C_{2}\theta^{2}n_{k}n$, neglecting the $k,q$ dependence of the rates $W_{k,q}^{\mp}$. The dominant heating mechanism is from pair creation in the bulk, $C_{\gamma}\gamma T(1-n_{k})(1-n)$. Then, we arrive at
@@
with $C=\pi^{2}C_{\gamma}^{\prime}/6C_{e}$. This expression holds in the weak-noise limit, $n^{\infty}N_{S}\sim{\gamma TN_{S}^{3}}/{\theta^{2}}\ll 1$, and we made use of the approximate quantization condition $k_{m}\approx\pi(1-m)/N_{S}$, which holds for $k_{m}\ll\pi$. These approximate rate equations, Eqs. (64,65), capture the crossover behaviour displayed in Fig. 5a.
@@
In Sec. III we argued that the fidelity of the steady state with the ground state approaches 1 as the MCP parameter $T\propto\beta$ is made large. In the presence of noise, there will be an optimal $T_{*}$ and $\beta_{*}$ that maximises the fidelity, since the noise strength also scales with $T$. We compare the log-fidelity per qubit for the MCP and the SCP vs. the noise ratio $\gamma/\theta^{2}$, in Fig. 5b. For each point, we optimize the parameter $\beta$, fixing $T=2\beta$ for simplicity (for the SCP this simply corresponds to optimizing $T$). We observe that, while at moderate noise values $\gamma/\theta^{2}\gtrsim 10^{-2}$, the SCP marginally outperforms MCP, for weak noise the MCP reaches lower fidelities in both phases, approaching fidelity 1 as the noiseless limit is approached. We conclude that the MCP is able to prepare target states with high fidelity even in the presence of weak noise, but noise is more destructive in systems with topological excitations such as the 1d AFM example.
V Cooling non-integrable systems
In this Section, we investigate the efficiency of the MCP for preparing ground states of non-integrable models. While these models are generally thermalizing, they typically have long-lived quasiparticles at low energy density. Therefore, the theory of dissipative cooling described in the previous sections for integrable systems, is expected to hold, once the system has been cooled to a state with sufficiently low energy density.
For a numerical test, we focus on the example of the TFIM with an additional longitudinal field,
$\hat{U}^{\prime}_{S}=e^{-i\hat{G}_{0}}e^{-i\hat{G}_{1}},$ (66)
Our parameter choice is in the Floquet regime, and not in the strict Trotter limit ($J,g_{x},g_{z}\ll O(1)$). Thus, to characterize the cooling efficiency, we consider effective Hamiltonians obtained in different orders of the Baker–Campbell–Hausdorff (BCH) expansion of Eq. (66), $\hat{U}_{S}^{\prime}\approx e^{-i\hat{H}_{l}}$, $l=0,1,\ldots$ (we will limit our consideration to $l=0,1$.) The first two orders of BCH formula give
where $l\in\{0,1\}$ denotes the BCH expansion order, and $\Omega$ is the corresponding ground state. The system is in the prethermal regime, due to the fact that $J,g,g_{X}\sim O(10^{-1})$, and is expected to cool towards the ground state of an effective Hamiltonian $\hat{H}_{l_{*}}$, with an optimal truncation order $l_{*}\geq 1$ of the BCH series. Thus, we expect the ground state of $\hat{H}_{1}$ to provide a better approximation than that of $\hat{H}_{0}$.
Figure 6: Cooling a non-integrable spin chain. (a) Evolution of energy, defined with respect to the ground state of approximate effective Hamiltonian $H_{0}$ (dashed lines) and $H_{1}$ (full lines), as a function of the number of resets $\nu$, for different system sizes. Corresponding energy values for the first excited states of $H_{1}$ are indicated by arrows, showing convergence of energy to a value well below the excitation gap. (b) Log-fidelity per qubit vs. number of resets. Model parameters are $h_{x}=h_{z}=0.15,J=0.1$ (PM phase), reset time $T=20=2\beta$. The bond dimension of simulations is $d=300$.
For the cooling protocol we focus on the MCP with finite density $N_{A}=N_{S}$. In contrast to the previous sections, where our aim was to develop a theory of cooling, here we generalise the protocol to cool more rapidly by using a large initial value of $\theta$, later approaching the weak coupling limit $\theta\ll 1$ where the previous analysis is expected to hold. To this end, we initialize the coupling at $\theta(0)\approx 0.4$ and progressively reduce to a final value $\theta_{\infty}\approx 0.083$. The value of the coupling is reduced as $\theta(t+1)=0.8\theta(t)$ whenever the energy of the system is reduced slower than a chosen rate, $0.010<E_{1}(t)-E_{1}(t-1)<$ 0.01. We found this choice to be efficient for the finite time simulations. Note that the function $f_{\tau}$ is still modulated within a cycle, see Eq. (29). We choose a reset time $T=20=2\beta$, for which $\Delta\beta\approx 2.6$, where $\Delta$ is the (many-body) energy gap of $\hat{H}_{1}$.
The tensor-network simulations are performed similarly to the integrable case, see Appendix E for details. The ground and first excited eigenstates of $\hat{H}_{0},\ \hat{H}_{1}$, and their corresponding energies, used in (68), are evaluated with a standard DMRG algorithm; the eigenstates are truncated to a bond dimension $d_{0}\approx\sqrt{d}$, where $d$ is the maximal bond dimension of the MPS performing the cooling protocol.
[…]
These results show that the MCP remains an effective tool for cooling non-integrable systems, which exhibit long-lived quasiparticles at a sufficiently low energy density. We expect that the performance of cooling protocols can be further improved, e.g. by adapting methods employed in variational quantum algorithms [68].

### v2

@@ III.3 Finite density of auxiliaries
To sum up, the analysis above demonstrates efficiency of MCP with a finite density of auxiliaries for removing quasiparticles after $O(\theta^{-2})$ cooling cycles, in the PM phase. In agreement with the arguments of Ref. [43], we also find that topological, domain-wall type quasiparticles are more challenging to remove. Nevertheless, in an idealized system without noise, our cooling protocol allows preparation of the ground state with very high fidelity, limited only by the choice of reset time $T$.
IV Cooling non-integrable systems
Having established the validity of our theory for the case of the TFIM, in this Section, we investigate the efficiency of the MCP for preparing ground states of non-integrable models. While these models are generally thermalizing, they typically have long-lived quasiparticles at low energy density. Therefore, the theory of dissipative cooling described in the previous sections for integrable systems, is expected to hold, once the system has been cooled to a state with sufficiently low energy density.
We study two examples of non-integrable systems using tensor network and quantum trajectory simulations: the TFIM in the paramagnetic regime with an additional longitudinal field [49], and an antiferromagnetic spin ladder with Heisenberg couplings between neighbouring sites [71]. In the longitudinal-field TFIM the physics is similar to the integrable case at low energies, except scattering now occurs between quasiparticles and multi-particle bound states emerge. Two-leg antiferromagnetic spin ladders model strongly-correlated low-dimensional magnetic systems [72, 73], with suggested connections to high-temperature superconductivity [50]. The low-energy spectrum features a band of spin-triplet excitations, as well as multi-particle bound states for intermediate coupling [71, 74]: thus, this relatively complex low-energy structure represents a good test for our cooling protocol.
IV.1 Longitudinal field TFIM
The longitudinal field TFIM is defined by the Floquet unitary
$\hat{U}^{\prime}_{S}=e^{-i\hat{G}_{0}}e^{-i\hat{G}_{1}},$ (54)
Our parameter choice is in the Floquet regime, and not in the strict Trotter limit ($J,g_{x},g_{z}\ll O(1)$). Thus, to characterize the cooling efficiency, we consider effective Hamiltonians obtained in the first two orders of the Baker–Campbell–Hausdorff (BCH) expansion of Eq. (54), $\hat{U}_{S}^{\prime}\approx e^{-i\hat{H}_{l}}$, $l=0,1$. The BCH formula gives
where $l\in\{0,1\}$ denotes the BCH expansion order, and $\Omega$ is the corresponding ground state. The system is in the prethermal regime, due to the fact that $J,g,g_{X}\sim O(0.1)$, and is expected to cool towards the ground state of an effective Hamiltonian $\hat{H}_{l_{*}}$, with an optimal truncation order $l_{*}\geq 1$ of the BCH series. Thus, we expect the ground state of $\hat{H}_{1}$ to provide a better approximation than that of $\hat{H}_{0}$.
Figure 5: Cooling a non-integrable spin chain. (a) Evolution of energy, defined with respect to the ground state energy $\varepsilon_{0}$ of approximate effective Hamiltonian $H_{0}$ (dashed lines) and $H_{1}$ (full lines), as a function of the number of resets $\nu$, for different system sizes. Corresponding energy values for the first excited states $\varepsilon_{1}$ of $H_{1}$ are indicated by arrows, showing convergence of energy to a value well below the excitation gap. (b) Log-fidelity per qubit vs. number of resets. Model parameters are $h_{x}=h_{z}=0.15,J=0.1$ (PM phase), reset time $T=\beta=20$. The bond dimension of simulations is $d=300$.
For the cooling protocol we focus on the MCP with finite density $N_{A}=N_{S}$. We take the coupling to auxiliaries to be the same as in Section III. In contrast to the previous sections, where our aim was to develop a theory of cooling, here we generalise the protocol to cool more rapidly by using a large initial value of $\theta$, later approaching the weak coupling limit $\theta\ll 1$ where the previous analysis is expected to hold. To this end, we initialize the coupling at $\theta(0)\approx 0.4$ and progressively reduce to a final value $\theta_{\infty}\approx 0.083$. The value of the coupling is reduced as $\theta(t+1)=0.8\theta(t)$ whenever the energy of the system is reduced slower than a chosen rate, $0.010<E_{1}(t)-E_{1}(t-1)<$ 0.01. We found this choice to be efficient for the finite time simulations. Note that the function $f_{\tau}$ is still modulated within a cycle, see Eq. (38). We choose a reset time $T=\beta=20$, for which $\Delta\beta\approx 5.2$, where $\Delta$ is the (many-body) energy gap of $\hat{H}_{1}$. The tensor-network simulations are performed similarly to the integrable case, see Appendix E for details.
[…]
IV.2 Antiferromagnetic Heisenberg ladder
As a second example, we consider a two-leg ladder geometry of $N_{S}=2\times L$ sites, with antiferromagnetic Heisenberg couplings between neighbouring sites within each chain and across the rungs [71], of strength $J$ and $J_{\perp}$ respectively. During the Floquet evolution, Heisenberg interactions are first applied to all odd-numbered bonds along the long (chain) axis of the ladder, followed by the even-numbered bonds, and finally across the rungs:
$\hat{U}_{S}^{\text{lad}}=e^{-i\hat{G}_{\perp}}e^{-i\hat{G}_{\text{even}}}e^{-i\hat{G}_{\text{odd}}},$ (57)
$\hat{G}_{\perp}=\frac{\pi J_{\perp}}{2}\sum_{\alpha,i=1}^{L}\hat{S}^{\alpha}_{1,i}\hat{S}^{\alpha}_{2,i},\;\ \hat{G}_{\text{even}}=\frac{\pi J}{2}\sum_{\sigma,\alpha,i=1}^{L/2-1}\hat{S}^{\alpha}_{\sigma,2i}\hat{S}^{\alpha}_{\sigma,2i+1},$
$\hat{G}_{\text{odd}}=\frac{\pi J}{2}\sum_{\sigma,\alpha,i=1}^{L/2}\hat{S}^{\alpha}_{\sigma,2i-1}\hat{S}^{\alpha}_{\sigma,2i},$
where $\alpha=x,y,z$ labels the spin component, $i$ runs along the long axis of the ladder, $\sigma=1,2$ labels the two sites belonging to each rung, and $\hat{S}^{\alpha}_{\sigma,i}$ are the spin-1/2 operators. Note the convention $\hat{S}^{x}=\hat{X}/2$ etc, and we have assumed that $L$ is even.
The non-perturbative regime $J\approx J_{\perp}$ is relevant to experiments [50]. In this parameter regime, the system is gapped and adiabatically connected to the point $J/J_{\perp}\approx 0$. The physics can be easily understood in this limit, where the ladder decouples into isolated antiferromagnetic rungs, each hosting an $S=0$ singlet-paired ground state with energy $-3J_{\perp}/4$, and a triplet of $S=1$ excitations with energy $J_{\perp}/4$ [71]. The corresponding quasiparticle creation operators (with annihilation operators identical due to Hermiticity) are specified by $\hat{Q}_{k}^{\alpha}=\frac{1}{\sqrt{N_{S}}}\sum_{j}e^{ikj}\hat{Q}_{j}^{\alpha}$, where
$\hat{Q}^{\alpha}_{i}=\hat{S}^{\alpha}_{1,i}-\hat{S}^{\alpha}_{2,i},$ (58)
and we have chosen the operators to anticommute with the parity operator exchanging the rung index $\sigma$. At lowest order in $J/J_{\perp}$, the quasiparticle band develops a finite dispersion as the excitations acquire a kinetic component along the ladder.
The adiabatic continuity argument suggests that the operators in Eq. (58) will continue to couple effectively to quasiparticles even in the regime $J/J_{\perp}\sim 1$. We therefore construct the cooling operators for the protocol to match the strong-coupling quasiparticle operators, taking the system-bath coupling as $\hat{V}_{j}^{\alpha}=\pi\hat{Q}^{\alpha}_{\mathcal{Q}_{j}}(\hat{\sigma}^{-}_{\mathcal{A}_{j}}+\sigma^{+}_{\mathcal{A}_{j}})$, with $\mathcal{Q}_{j}$ ($\mathcal{A}_{j}$) labelling the system (auxiliary) position associated with $j$-th coupling. Each auxiliary qubit couples with a single choice of $\alpha$ during each cooling cycle (in the numerics we randomly assign the couplings at the start of each cycle).
We test the effectiveness of the above cooling protocol, using the MCP with a finite density of auxiliary qubits. For the simulations we employ the quantum trajectory method, using Google’s qsim simulator [75]. Each trajectory represents an exact evolution of the system’s wavefunction, conditioned on a stochastic sequence of events associated with the auxiliary reset. The initial state for each trajectory is a random computational basis state, uniformly sampled from an infinite-temperature ensemble. Similarly to the longitudinal-field TFIM, we measure the system energy with respect to the first two orders of the BCH expansion. We also track the system’s average magnetization along the $\hat{z}$-axis, $\langle\hat{S}_{i}^{z}\rangle$, which is expected to be zero in the ground state due to the spin-symmetry of the Heisenberg terms.
The results are shown in Fig. 6, for a ladder of $2\times 10$ qubits, coupled to 6 auxiliary qubits. The couplings $\mathcal{A}_{j}(\mathcal{Q}_{j})$ are set randomly at the start of each cycle. We take $J=J_{\perp}$ = 0.2, which is in the non-perturbative regime discussed above. The coupling strength is $\theta=0.15$ and reset time $T=\beta=30$. For these parameters the excitation gap of $\hat{H}_{1}$ is $\Delta=0.19$ and $\Delta\beta\approx 5.8$. We observe the measured energies converging closely toward the ground state energy, with the final $H_{1}$ energy being slightly closer to the ground state than for $H_{0}$. At late times, the magnitude of the energy fluctuations falls below the excitation gap, indicative of a large overlap between the steady state and the system ground state. The measurements of the average magnetisation, displayed in the inset, show that the system approaches a magnetically disordered state on even faster timescales. We checked several other values of $J<J_{\perp}$ and found that the ground state convergence is robust. In Appendix F, we provide additional data showing that the final energy density and time to reach the steady state is independent of system size, similarly to the TFIM paramagnet, assuming a finite density of auxiliary qubits. This agrees with our expectation that systems adiabatically connected to a fixed point with local excitations, can be cooled in constant time.
Together, the results for the longitudinal-field Ising model and Heisenberg spin ladder show that the MCP remains an effective tool for cooling non-integrable systems.
Figure 6: Cooling of a Heisenberg spin ladder. The mean energy for approximate effective Hamiltonians $H_{0}$ (dashed) and $H_{1}$ (solid) is shown, relative to the ground state energy, vs. the number of resets $\nu$. The average is calculated over 100 trajectories, with the blue shaded area showing the standard deviation of the $H_{1}$ data. The excitation gap of $H_{1}$ is marked by the red line. Inset: Evolution of average magnetisation vs. number of resets. We use $J=J_{\perp}=0.2$, $\theta=0.15$, and the MCP with $T=\beta=30$.
V Noise effects
@@
The results of the previous sections demonstrate the efficiency of our cooling protocol for preparing correlated ground states, assuming the reset time $T$ can be made large, which is the case in the absence of noise. We now turn to the effects of unwanted decoherence. Each experimental platform comes with its own sources of noise, and our goal will be to understand the limitations it imposes on the cooling efficiency. We will focus on the single-qubit decay and dephasing processes, relevant in systems of superconducting qubits. The kinetic equation can be generalized to include transitions mediated by weak noise. Using this modified kinetic theory, we study the 1d TFIM model and analyze the dependence of the steady state quasiparticle population on the noise strength.
V.1 Rate equation in the presence of noise
@@
Considering continuous time evolution, single-qubit decoherence processes can be conveniently described within the Lindblad formalism [52]. The evolution of the system’s density matrix $\hat{\rho}_{S}$ is given by (for the moment, we ignore the presence of auxiliary qubits):
@@
and following Eqs. (16-25) we obtain the noisy rate equation, which includes cooling transitions $R_{\vec{\alpha}\vec{\beta}}$ and and an additional contribution $R_{\vec{\alpha}\vec{\beta}}^{\gamma}$ to transition rates from the noise terms:
@@
$R_{\vec{\alpha}\vec{\beta}}$ $=\theta^{2}\sum_{a}|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}|\bra{\vec{\beta}}\hat{A}^{a}\ket{\vec{\alpha}}|^{2},$ (65)
@@
V.2 Application: cooling the transverse-field Ising model
Figure 7: Performance of the cooling algorithm under decoherence. (a) Quasiparticle density in the steady state, as a function of noise ratio $\gamma/\theta^{2}$, in PM (red) and AFM (blue) phases, for different system sizes $N_{S}$ from 10 to 80. We use MCP with fixed $T=\beta=6\pi/\Delta$. At weak noise, PM quasiparticle density scales linearly with $\gamma$, while AFM displays a crossover from $n\propto\gamma$ to $n\propto\sqrt{\gamma}$ behaviour as noise strength is reduced. (b) Log-fidelity between steady state and quasiparticle vacuum, for MCP (solid lines) and SCP (dashed lines). We fix $T=\beta$ and optimize over parameter $\beta$ for each protocol.
Next, we apply the noisy kinetic equation to cooling of the transverse-field Ising model of Section III in the presence of decay to the qubit ground state, and dephasing noise. This corresponds to $\mu=d,\phi$, with jump operators
@@ V.2 Application: cooling the transverse-field Ising model
We assume homogeneous decoherence rates for simplicity, $\gamma_{j\phi}=\gamma_{\phi}$, $\gamma_{jd}=\gamma_{d}$, and rescale the dephasing rate, $\gamma_{\phi}\to\gamma_{\phi}/2$, following standard conventions in the literature on open quantum system. Non-uniform decoherence rates can be further included in the rate equation without complication (we refer to [76] for a characterisation of site-dependent noise rates for the Google Quantum Processor).
@@
The matrix elements satisfy the selection rules discussed in Sec. III. Note that while the matrix elements for cooling and decay processes are identical here, the spectral filtering provided by the filter function distinguishes the two processes. We now summarise the main behaviour of the noisy rate equation in the paramagnetic and antiferromagnetic phases of the TFIM, with a more detailed discussion of matrix elements for $\hat{Z}_{j}$ and $\hat{\sigma}^{\pm}_{j}$ operators given in Appendix D. In the PM phase, single particle loss and gain processes occur throughout the chain (cooling and decay contributions); quasiparticle pair creation/annihilation and number-conserving scattering events occur due to dephasing. In the AFM phase, on account of the domain-wall nature of excitations, only two-particle scattering and pair creation/annihilation processes can occur in the bulk, and they are induced by cooling, decay and dephasing. Single particle transitions require an interaction with the Majorana edge mode, and are localized to the chain edge.
We have compared the predictions of the noisy kinetic equation and MPS simulations in the presence of noise, finding again a good agreement in the parameter regimes accessible to MPS (moderate value of noise strength, or moderate times at a very weak noise). We refer the reader to Appendix G for a comparison of the steady state occupations, also for different coupling strengths. The rate equation, however, allows us to study system sizes and parameter regimes far beyond the reach of matrix product state simulations  [^5] The steady state can be efficiently found by root-finding algorithms, after the cooling and noise-induced rates have been calculated via Wick’s theorem. In practice, the system reaches its steady state quickly and we evolve the rate equation in time until the quasiparticle occupations have effectively converged..
In Fig. 7a, we illustrate the quasiparticle density in the steady state, obtained from the kinetic equation, as a function of the noise strength. For simplicity, we choose $\gamma_{d}=\gamma_{\phi}=\gamma$ and fix auxiliary density $N_{A}=N_{S}$. We focus on the MCP with $T=\beta=6\pi/\Delta$. We observe a qualitatively different behavior in the two phases, in line with above results on noiseless cooling. In the PM phase, the steady state quasiparticle density decays as a power-law function of noise, $n^{\infty}_{\rm PM}\propto\gamma^{a}$, $a\approx 1$, and is independent of the system size for $N_{S}\gtrsim 10$. In the AFM phase, in contrast, the density displays a power-law scaling which depends on system size, $n^{\infty}_{\rm AFM}\propto\gamma^{b(N_{S})}$. For small systems and weak noise, we observe similar scaling to the PM, with $b\approx 1$; as either the system size or noise strength is increased, we find a crossover to a different power $b\approx 1/2$.
@@
where $C_{1},C_{\gamma}$ are constants of order one. We assumed that MCP is in the regime where heating processes due to coupling to auxiliaries are strongly suppressed, and neglected the dependence of $C_{1},C_{\gamma}$ on $k$. For the noise we neglect terms $O(\gamma Tn)$, since we are interested in the regime of weak noise and low density. Then, the expected steady state quasiparticle density is given by
@@
(strictly speaking, this expression holds for small quasimomenta, $k\ll 1$, but we will see below that the steady state quasiparticle density is determined mostly by this quasimomentum range). The processes removing quasiparticle pairs can be modeled by a term $-C_{2}\theta^{2}n_{k}n$, neglecting the $k,q$ dependence of the rates $W_{k,q}^{\mp}$. The dominant heating mechanism is from pair creation in the bulk. Then, we arrive at
@@
with $C=\pi^{2}C_{\gamma}^{\prime}/(6C_{e})$. This expression holds in the weak-noise limit, $n^{\infty}N_{S}\sim{\gamma TN_{S}^{3}}/{\theta^{2}}\ll 1$, and we made use of the approximate quantization condition $k_{m}\approx\pi(1-m)/N_{S}$, which holds for $k_{m}\ll\pi$. These approximate rate equations, Eqs. (77,78), capture the crossover behaviour displayed in Fig. 7a.
@@
In Sec. III we argued that the fidelity of the steady state with the ground state approaches 1 as the MCP parameter $T\propto\beta$ is made large. In the presence of noise, there will be an optimal $T_{*}$ and $\beta_{*}$ that maximise the fidelity, since the noise strength also scales with $T$. We compare the log-fidelity per qubit for the MCP and the SCP vs. the noise ratio $\gamma/\theta^{2}$, in Fig. 7b. For each point, we optimize the parameter $\beta$, fixing $T=\beta$ for simplicity (for the SCP this simply corresponds to optimizing $T$). We observe that, while at moderate noise values $\gamma/\theta^{2}\gtrsim 10^{-2}$, the SCP marginally outperforms MCP, for weak noise the MCP reaches higher fidelities in both phases, approaching fidelity 1 as the noiseless limit is approached. We conclude that the MCP is able to prepare target states with high fidelity even in the presence of weak noise, but noise is more destructive in systems with topological excitations such as the 1d AFM example.